### v1

@@ 3.4 Algorithm: Matching Candidate Messages
During iteration, we need to decide whether $m_{S}$ can be mapped to a received message. We do so by comparing $\psi(m_{S},r)$ to the (previous) message anonymity set $\psi(m_{S}^{\prime},r)$ of the previous message $m_{S}^{\prime}$. $m_{S}$ can mapped if and only if there is at least one valid candidate message. The number of candidates (cand) is determined as the sum of leftover unused candidate messages from the previous anonymity set $\psi(m_{S}^{\prime},r)$ that are also present in $\psi(m_{S},r)$, which we call overlap, plus the number of new messages $|\psi(m_{S},r)\setminus\psi(m_{S}^{\prime},r)|$, which we call new. If cand > 0, then $m_{S}$ can be mapped and the next message continues with cand - 1. If for all messages, the number of candidates is non-zero, then a suitable assignment $\sigma$ exists. Otherwise, this is not the case.
3.5 Implementation: ppcalc
@@
We implement progressive pruning and make it available as an open-source project called ppcalc  [^1] https://github.com/cdoepmann/ppcalc, implemented in the Rust language. ppcalc consumes network traces consisting of messages between senders and receivers, and their timestamps. It outputs the anonymity set of each stream. It processes plain CSV data and can thus easily be reused for other research questions. It supports the calculation of sender anonymity as well as receiver anonymity. During the implementation of ppcalc, great care was taken to achieve high scalability that enables ppcalc to process communication traces not only from relatively small, synthetic scenarios as considered in this section, but also traces from real-world networks like Tor, at scale. For the experiments we will later present in Section 6, ppcalc successfully managed to process network traces of approximately 720 GB within 74 hours on a 12 CPU cores machine.
4 Understanding Anonymity Over Time
Progressive pruning, as presented in Section 3.2, explicitly focuses on the evaluation of anonymity within stream-based communication. That is, it takes into account how the anonymity set changes over the course of a data transfer that consists of a sequence of messages. In order to understand the implications of this measurement approach, we therefore analyze how anonymity changes with an ongoing stream of messages. This can also be understood as a development over time as the single messages are generally sent sequentially. Note that progressive pruning does not require any specific temporal characteristics but works the same if all messages in a stream are sent simultaneously.
For this first analysis, we rely purely on stochastical methods in order to learn about the time behavior that is inherent to our model. In order to do so, we first consider the following very simple model of communication over an anonymous communication network: For now, we assume that all users send their messages at the same rate. This assumption allows us to discretize the system into synchronized rounds, assuming that one round corresponds to the transmission of exactly one message by each user that is active during that round. Although this is a simplification, we argue that is still close to the overall model behavior due to its window semantics. Accordingly, the window size that defines which messages belong to the anonymity set corresponds to one round. Making use of the terminology from Section 3.2, messages that are sent in the same round form a common message-message anonymity set. More intuitively, senders belong to the same anonymity set if they send data at the same time (within the same window/round). This matches our intuitive understanding of traffic analysis, as users sending messages at clearly different times are easy to tell apart. For simplicity, we further assume that all senders transfer the same number $L$ of messages, in $L$ consecutive rounds.
Let us consider a single sender $s$ within a busy anonymity system, starting to send its messages at round 00. Then, the transfer will last $L$ rounds. Moreover, let $N$ be the overall number of (independent) senders active in the system at this first round. For each round $t$, we want to find the anonymity set size $|\Psi_{SR}[s]|$, which we abbreviate as $A_{t}$ here. Obviously, $A_{0}=N$. Since our sender $s$ starts sending while the system is already in use by other senders, those senders will finish sending their messages at different times. Let $E_{i}$ be the time when sender $i$’s transfer ends (more precisely, the first round in which sender $i$ does not send a message anymore). Then, $E_{s}=L$ because $s$ sends for $L$ rounds and starts at round 00.
In this scenario, the anonymity set will only shrink when other senders stop sending at some time. Note that users starting to send later do not have an impact on the anonymity set because they have not been candidates before and thus cannot be valid anonymity set members later. The anonymity set cannot grow over time because it corresponds to the intersection of observations over time. The anonymity set size is therefore determined by the number of senders that have stopped sending. Let $\bar{A}_{t}$ be the number of senders that have already finished their transfer before round $t$. Then the anonymity set size is $A_{t}=N-\bar{A}_{t}$, and $\bar{A}_{0}=0$.
Without loss of generality, let us assume that $E_{s}$ corresponds to $E_{0}$, simplifying the following formulations. Since all senders transfer the same number $L$ of messages, and are independent from each other, we can calculate the probabilities for certain end times as follows:
$P(E_{i}=1)=\dots=P(E_{i}=L)=\frac{1}{L}~{}~{}~{}\text{for $0<i\leq N$}$
Consequently, we can accumulate the probabilities as follows (note that the events $\{E_{i}=t\}$ are all disjoint):
$P(E_{i}\leq 0)=0\qquad P(E_{i}\leq L)=1\qquad P(E_{i}\leq t)=t\cdot\frac{1}{L}$
The relationship between $\bar{A}_{t}$ and the different $E_{i}$ is as follows: For a fixed round $t$, $\bar{A}_{t}=t$ iff. $E_{i}\leq t$ for $x$ different $i$, and $E_{i}>t$ for the remaining $(N-x)$ $E_{i}$. Due to the fact that the events $\{E_{i}\leq t\}$ are all independent of each other, we can model each of them as a Bernoulli experiment with success probability $p=t\cdot\frac{1}{L}$. For each round $t$, $\bar{A}_{t}$ is therefore binomially distributed with $n=N$ and $p=t\cdot\frac{1}{L}$:
$P(\bar{A}_{t}=x)=\binom{n}{x}\cdot p^{x}\cdot(1-p)^{n-x}$
Consequently, the anonymity set size is distributed as follows:
$P(A_{t}=x)=P(\bar{A}_{t}=N-x)=\binom{N}{N-x}\cdot p^{N-x}\cdot(1-p)^{x}\qquad\square$
We can draw interesting insights about the anonymity set size as modeled by our progressive pruning. In particular, if we calculate the expected value for the anonymity set size in round $t$, we get:
$E[A_{t}]=E[N-\bar{A}_{t}]=N-E[\bar{A}_{t}]=N-n\cdot p=N-N\cdot\left(t\cdot\frac{1}{L}\right)$
As a consequence, we can see that the expected anonymity set size decreases linearly/anti-proportionally over time. Put differently, the longer the transfer lasts, the weaker the achieved anonymity gets. This observation clearly underlines the significance of the data stream length: The shorter the communication is, the more likely it is to hide among other senders. Also note that the simple system setup described here (same-length senders with independent start times) is equivalent in principle to another relevant scenario to consider in which all senders start sending at the same time but have independently selected transfer lengths.
@@ 5 Anonymity Factors for Streams
{subcaptionblock}
.3 Refer to caption {subcaptionblock}.3 Refer to caption {subcaptionblock}.3 Refer to caption
Figure 3:
Figure 4:
Figure 5:
Figure 6: Influence of different stream properties on stream anonymity set size.
@@
Our general approach is as follows: We simulate communication between abstract senders and receivers with different scenario parameters. This simulation of communication streams is carried out on a high abstraction level—the communication properties are defined only by general parameters like stream length, latency, and number of senders and receivers. We then apply progressive pruning and use the calculated anonymity set sizes as estimates for anonymity. By altering the scenario parameters, we can evaluate the relative influence of each parameter on anonymity. Without loss of generality, we focus on sender anonymity here. Therefore, each sender sends all of their messages to one receiver, and the anonymity set of a sender corresponds to the set of candidate receivers as observed by an attacker.
@@
As our base scenario, we choose 2,000 senders, starting their transfer within a time span of 60 seconds. This number is loosely inspired by the typical 3 million daily users in Tor. Each sender transmits 2,314 KB of data, corresponding to the average size of a website. [^2] https://almanac.httparchive.org/en/2022/page-weight#request-bytes Senders transmit at a data rate of 10 Mbit/s, sending to a total of 400 receivers, chosen uniformly at random. The network latency is set to 210 ms, a typical value found in the real Tor network. [^3] https://metrics.torproject.org/onionperf-latencies.csv?start=2023-05-05&end=2023-08-03&server=public The analysis window that bounds network latency for anonymity calculation in progressive pruning is set to ±$\pm 100$ ms around this value. For each of the parameters, we carry out a series of simulations in which we only vary this parameter.
Our experiments reveal and verify the following relationships: Firstly, the length of a stream, in relation to other streams in the network, is of crucial importance for the achieved anonymity (cf. Figure 6). While short streams can achieve large anonymity sets, longer streams serve as cover traffic for the whole network but lose their own anonymity. This underlines the effectiveness of cover traffic, either implicitly through a large user base, or explicitly using designated cover traffic. Secondly, a low number of receivers compared to the number of senders is beneficial for anonymity because it helps streams to blend in with each other (Figure 6). And thirdly, homogeneity of streams is also beneficial for anonymity. (Figure 6).
@@
While these results come as no surprise, they express the soundness and utility of predictive pruning to analyze anonymity properties of streams, also verifying these relationships based on a well-defined stream model.
@@ 6 Case Study: Comparing Tor Instances
In the previous section, we demonstrated the behavior of progressive pruning for the analysis of stream-based communication, and inferred general factors that are relevant for the achieved anonymity. The considered scenarios, however, were synthetic examples suitable primarily for investigating and verifying our metric as well as general, abstract trends. In contrast, we now go further and aim to apply our analysis approach to the most widely used anonymity network as of today—the Tor network. By doing so, we want to give a practical example of how progressive pruning can be used to research the anonymity properties of large-scale anonymity systems providing stream communication.
As laid out in Section 3, progressive pruning is not suitable for quantifying anonymity as an absolute value, but rather as a relative measure for comparing anonymity. We therefore refrain from quantifying “the anonymity level of Tor”. The quantification can, however, be used to compare different instances of the network, hinting at deviations in achieved anonymity. We use this approach to investigate the implications of one of the future changes Tor will face, namely network growth.
Over time, the Tor network has grown significantly and it is expected to continue this trend in the future [DBLP:conf/IEEEares/DopmannT23]. While this development indicates the increasing significance and the growing user base of Tor, it does not come without challenges. Upcoming issues should be spotted early, so they can be taken into consideration when designing future network iterations. In [DBLP:conf/IEEEares/DopmannT23], the authors introduced a methodology to investigate upscaled instances of the Tor network, with regard to different topological properties, also finding initial evidence that purely horizontal growth of the Tor network can lead to suboptimal performance characteristics. We now aim to leverage our progressive pruning methodology to analyze the scaling impact on anonymity provided by the network.
@@ 6.1 Methodology
Our general approach is to estimate the achieved anonymity based on simulations of the network. For this, we first simulate the Tor network in order to obtain a set of circuits that were created, along with traces of the exchanged messages (cells) and their timestamps. We then compute the anonymity level according to progressive pruning. This process is first carried out for the original (unscaled) Tor network. Using torsynth [DBLP:conf/IEEEares/DopmannT23], we then generate two synthetic but representative Tor networks that were scaled either vertically or horizontally by +100%. We also evaluate these networks using the aforementioned workflow and then compare the achieved anonymity values. In order to avoid sampling artifacts, we refrain from down-sampling the networks for simulation. Instead, we carry out all experiments at scale, which constitutes a nontrivial engineering challenge to scalability of the involved software components. We base our experiments on the network consensus from 2023-10-01 at 00:00, and simulate one hour of full network activity.
@@
Given a specific snapshot of the Tor network (consensus), the network simulation step needs to generate a trace of network activity within the Tor network. In particular, this involves generating circuits and using them to transport data cells as would be the case in the real network.
@@
There are different approaches towards simulating the Tor network, trading between computational effort and realism. The most realistic simulations can be carried out using the shadow simulator which is able to run the unmodified Tor client. However, running such simulations at scale 1:1 (or even 2:1 in the case of up-scaled networks) was infeasible. While shadow enables a very fine-grained simulation to investigate performance and other networking effects, we are particularly interested in how circuits and cells are distributed across the network because these are the factors for anonymity as considered by our metric. We therefore choose a more abstract simulation approach, similar to the Tor path simulator (TorPS [johnson2013usersrouted]). TorPS simulates circuit generation by Tor users over time and has been used before to evaluate how anonymity changes when users construct many consecutive circuits. However, TorPS suffers from severe scalability issues. Furthermore, it does not simulate user behavior at the packet level and thus cannot be used to generate the network traces we require.
@@
We therefore propose a novel simulator software which we call Tor Flow Simulator (TorFS) and make available as an open source project [^4] https://github.com/cdoepmann/torfs, TorFS inherits much of TorPS’s functionality. However, it is implemented in the Rust programming language with a focus on scalability, and adds the missing pieces needed to simulate Tor user behavior at the packet level. Most importantly, it applies the modeled circuit selection rules as well as the users’ circuit management. It also supersedes TorPS’s simplistic user model, replacing it with the state-of-the-art model introduced in [DBLP:conf/ccs/JansenTH18] and used by other publications in the field since then [DBLP:journals/popets/PahlAKE23, DBLP:conf/ccs/RochetWJMP20, DBLP:conf/uss/JansenVS19]. As in [DBLP:conf/ccs/JansenTH18], user behavior is now modeled as a probabilistic process (an exponential distribution) with parameters derived from real-world measurements on the Tor network. Moreover, while TorPS simulates only the generation and management of circuits, TorFS also simulates the circuits themselves to generate network traces of the individual messages (cells). For this, it again makes use of the probabilistic model from [DBLP:conf/ccs/JansenTH18]. Consequently, TorFS uses the stream model as well as the packet model from [DBLP:conf/ccs/JansenTH18] which are Markov chains to mimic real traffic behavior as observed on the real Tor network. At the same time, TorFS does not constitute a full-fledged network simulator like shadow or ns-3. It abstracts from networking effects in that cells entering the network are output after a constant latency at their destination. This makes it feasible to simulate the network at scale. Moreover, we argue that this is in line with our approach of focusing on the inherent anonymity effects within the Tor network as influenced by its topology.
@@
TorFS outputs message traces ready for analysis with ppcalc. We only focus on traffic that is directed from the Tor exit to the client because this constitutes the majority of Tor traffic. We apply the following mapping from simulated Tor behavior to the abstract message model from Section 3: Each cell that occurs in the simulation corresponds to a message in the abstract model. Multiple cells that form a data stream in the simulation also constitute a single stream of messages in the sense of our abstract model. The abstract sender of these messages is the according simulated exit. For each stream, we define an abstract receiver as the entity that receives all of this stream’s messages. This way, the stream-based anonymity can be evaluated independently for each stream in the network.
Anonymity Computation:
@@
After generating message traces of the Tor network using TorFS, we analyze these using progressive pruning in order to estimate the anonymity level. For this task, we again used ppcalc (cf. Section 3.5). As laid out before, the anonymity estimate can be calculated in either one of two possible “directions”. In our current experiment, we have a fixed (relatively small) set of exit relays that are sending data to many clients. Due to this 1:n relationship, it makes sense to calculate anonymity from the receiver perspective. That is, for each receiver (client), we calculate the set of potential senders (exits) from which their data may originate. Analyzing anonymity in the reverse direction (the set of potential receivers per sender) is not useful in this scenario due to the fact that exits are sending to many clients at the same time.
@@ 6.2 Results and Discussion
{subcaptionblock}
.45 Refer to caption {subcaptionblock}.45 Refer to caption
Figure 7: Orig. (Refer to caption), vert. (Refer to caption), horiz. (Refer to caption)
Figure 8: Original
{subcaptionblock}
.45 Refer to caption {subcaptionblock}.45 Refer to caption
Figure 9: Vertically scaled
Figure 10: Horizontally scaled
Figure 11: (a) Overall anonymity distribution of streams per scenario, (b)-(d) Anonymity distribution of streams, grouped by stream length per scenario (from percentiles 0-10/black, to percentiles 90-100/orange).
@@
Our analysis is based on the streams’ anonymity set sizes and their distribution across the network. Table 1 displays fundamental statistics about the anonymity in each simulated network. As a first observation, we can see that, on average, the anonymity set sizes are relatively large. In the case of the unmodified Tor network, e.g., streams still remain with an average of more than 1,200 sender candidates (out of a total number of approximately 2,400 exits). It is noteworthy that, with regard to the simple window-based attack progressive pruning mimics, anonymity of the vast majority of streams is not trivially broken. This observation is supported by the fact that—looking again at the original Tor network—out of 26 million streams, less than 1,000 are fully de-anonymized (resulting in an anonymity set size of 1).
@@
In order to better understand how anonymity is distributed within the network, we consider CDF plots of the achieved anonymity set sizes, displayed in Figure 11. Looking at the data from the original Tor network first, it becomes immediately apparent that anonymity is distributed very unevenly between the streams. Note that, if all streams had the same anonymity set size, the CDF plots would exhibit a single vertical line. In contrast, the simulation results show that there are many streams with very little anonymity remaining, just like there are many streams with rather strong anonymity remaining and many streams in between. While a result like this was to be expected due to the fact that Tor has no means of ensuring a certain level of anonymity, it still shows a surprisingly clear weakness of the system because users cannot reasonably expect a certain level of anonymity, as far as the stream-based anonymity degradation over time is concerned.
@@
Having seen in Section 5 that a stream’s anonymity is heavily influenced by its length, we next investigate the relationship between stream lengths and anonymity in the simulated networks. Figure 11 contains CDF plots of the achieved anonymity set sizes again, but subdivides the streams into quantiles based on their length (number of cells). The resulting anonymity distributions clearly point out that stream length is again a decisive factor for the resulting anonymity. Long-lasting streams are massively disadvantaged and generally achieve only little anonymity. In contrast, short streams are much harder to de-anonymize and have anonymity sets that are multiple times larger. These results again support our observations from Section 5: Long streams suffer from reduced anonymity but act as cover traffic providing stronger anonymity for shorter streams.
Anonymity in Scaled Networks
@@
We leverage our simulation and analysis workflow to evaluate not only the Tor network as is today but also variations of the network that could result from future network growth of different characteristics. Our choice to evaluate scaled Tor networks is motivated by two goals: Firstly, it allows us to research the influence of network topology on the achieved anonymity. And secondly, understanding the trade-offs of different topological properties may provide important insights on how to shape and prepare for future network evolution. In addition to a snapshot of today’s Tor network, we therefore carried out simulations of the Tor network being scaled by +100% vertically as well as by +100% horizontally. By “vertical” growth we mean that the bandwidth capacity of each relay is increased accordingly. Likewise, “horizontal” growth is achieved by adding new, additional relays to the existing network. The scaling methodology of torsynth is detailed in [DBLP:conf/IEEEares/DopmannT23]. For simulations, we also grow the number of active users on the network by +100% so the networks differ only in their topology, and not in the amount of available resources per user.
@@
We first consider the case of horizontal growth. Table 1 and Figure 11 show that increasing the number of relays has a strong impact on anonymity as defined by progressive pruning. In fact, with double the number of relays, the average anonymity size is also exactly doubled. This trend is intuitively reasonable because doubling the number of relays also implies doubling the number of available exits which are the potential anonymity set candidates. However, it is interesting to see that this increase of average anonymity set size appears to be exactly proportional to the number of exit relays. This is somewhat surprising because, essentially, the number of exits only directly defines the maximum anonymity set size which wouldn’t necessarily have to correspond linearly to the average anonymity set size. It is noteworthy that the standard deviation also doubles exactly.
@@
All in all, we can conclude from the simulations of the Tor network which we carried out that there are two fundamental factors for anonymity in the network: Firstly, an increased number of relays also increases the number of (topological) candidates for the anonymity set. And secondly, apart from the topological factor, the mere increase in network load also improves anonymity, albeit much less effectively.

### v2

@@ 3.5 Algorithm for Matching Candidate Messages
During iteration, we need to decide whether $m_{S}$ can be mapped to a received message. We do so by comparing $\psi(m_{S},r)$ to the (previous) message anonymity set $\psi(m_{S}^{\prime},r)$ of the previous message $m_{S}^{\prime}$. $m_{S}$ can be mapped if and only if there is at least one valid candidate message. The number of candidates (cand) is determined as the sum of leftover unused candidate messages from the previous anonymity set $\psi(m_{S}^{\prime},r)$ that are also present in $\psi(m_{S},r)$, which we call overlap, plus the number of new messages $|\psi(m_{S},r)\setminus\psi(m_{S}^{\prime},r)|$, which we call new. If cand > 0, then $m_{S}$ can be mapped and the next message continues with cand - 1. If for all messages, the number of candidates is non-zero, then a suitable assignment $\sigma$ exists. Otherwise, this is not the case.
3.6 Implementation
@@
We implement progressive pruning and make it available as an open-source project called ppcalc  [^1] https://github.com/cdoepmann/ppcalc, implemented in Rust. ppcalc consumes network traces consisting of messages between senders and receivers, and their timestamps. It outputs the anonymity set of each stream. It processes plain CSV data and can thus easily be reused for other research questions. It supports the calculation of sender anonymity as well as receiver anonymity. During the implementation of ppcalc, great care was taken to achieve high scalability that enables ppcalc to process communication traces not only from relatively small, synthetic scenarios but also traces from real-world networks like Tor, at scale. For the experiments we will later present in Section 5, ppcalc successfully managed to process network traces of approximately 720 GB within 74 hours on a 12 CPU cores machine.
@@ 4 Anonymity Factors for Streams
[…]
Figure 3: Influence of different stream properties on the anonymity set size.
@@
To this end, we simulate communication between senders and receivers with different scenario parameters such as stream length, latency, and number of senders and receivers. We then apply progressive pruning and use the calculated anonymity set sizes as estimates for anonymity. By altering the scenario parameters, we can evaluate the relative influence of each parameter on anonymity. Without loss of generality, we focus on sender anonymity here. Therefore, each sender sends all of their messages to one receiver, and the anonymity set of a sender corresponds to the set of candidate receivers as observed by an attacker.
@@
As our base scenario, we choose 2,000 senders, starting their transfer within a time span of 60 seconds. This number is loosely inspired by the typical 3 million daily users in Tor. Each sender transmits 2,314 KB of data, corresponding to the average size of a website. [^2] https://almanac.httparchive.org/en/2022/page-weight#request-bytes Senders transmit at a data rate of 10 Mbit/s, sending to a total of 400 receivers, chosen uniformly at random. The network latency is set to 210 ms, a typical value found in the real Tor network. [^3] https://metrics.torproject.org/onionperf-latencies.csv?start=2023-05-05&end=2023-08-03&server=public The analysis window that bounds network latency for anonymity calculation in progressive pruning is set to ±$\pm 100$ ms around this value. For each of the parameters, we carry out a series of simulations in which we only vary this parameter. In the following, we describe our results, revealing and verifying several relationships.
Stream Length
Firstly, the length of a stream, in relation to other streams in the network, is of crucial importance for the achieved anonymity (Figure 3). The results reveal that anonymity sets increase with growing stream length variation. At first sight, this is counterintuitive because more heterogeneous stream lengths may give an attacker a way to distinguish between streams. However, what can be observed here is the effect of cover traffic: Longer streams act as cover traffic for shorter streams, enhancing their anonymity while sacrificing their own. Due to the fact that a single long stream can act as cover traffic for several shorter streams, the average anonymity is increased. This underlines the effectiveness of cover traffic, either implicitly through a large user base, or explicitly using designated cover traffic.
Network Topology
We analyze the impact of network topology by varying the ratio between senders and receivers. With fewer receivers, each receiver handles the streams of more senders, leading to a higher-density network.
We find that a low number of receivers compared to the number of senders is beneficial for anonymity because it helps streams to blend in with each other (Figure 3). This is in line with our expectation because a larger number of receivers leads to an increasingly sparse network, in which there are fewer candidates an attacker has to differentiate. However, a too low number of receivers is problematic for our notion of anonymity as well because the total number of receivers constitutes an inherent upper bound for the anonymity set size.
Network Latency
Lastly, we consider the case of varying network lateny (jitter). The results in Figure 3 indicate a small increase in anonymity with growing network jitter. There are two main reasons for this: Firstly, more network jitter generally makes the network less predictable for an attacker. And secondly, he introduction of jitter for the single messages effectively results in the streams spanning an increased period of time during transmission. If streams are active for a longer time, they will be active with more other streams simultaneously, leading to an increased anonymity set.
@@
All in all, these results come as no surprise, but express the soundness and utility of predictive pruning to analyze anonymity properties of streams, also verifying these relationships based on a well-defined stream model.
@@ 5 Case Study: Comparing Tor Instances
So far, we demonstrated how progressive pruning can be used for analyzing stream-based communication and identifying factors relevant to anonymity. The considered scenarios, however, were primarily synthetic examples to verify our methodology and observe trends. Now, we apply this approach to the widely used Tor network, providing a practical example of how progressive pruning can assess the anonymity properties of large-scale systems.
As discussed, progressive pruning quantifies anonymity relatively. Instead of defining “Tor’s anonymity level,” we compare network instances to identify deviations in anonymity, focusing on the implications of network growth.
The Tor network has grown significantly over time and it is expected to continue this trend. This growth, however, presents challenges for future network designs. Previous work [DBLP:conf/IEEEares/DopmannT23] analyzed scaled Tor instances, revealing that purely horizontal growth may harm performance. Here, we leverage progressive pruning to explore the impact of scaling on network anonymity, which is necessary as adversaries become stronger.
@@ 5.1 Methodology
Our general approach is to quantify anonymity sets based on message traces obtained from network simulations. Specifically, we simulate the Tor network to obtain a set of circuits, along with traces of exchanged messages (cells) and their timestamps. We then compute anonymity sets over time using progressive pruning. This process is first applied to the original (unscaled) Tor network. In addition, we generate two synthetic but representative Tor networks scaled either vertically or horizontally by +100% using torsynth [DBLP:conf/IEEEares/DopmannT23]. We conduct network simulations on these scaled networks and apply progressive pruning to the resulting traces to quantify anonymity sets as well.
In order to avoid sampling artifacts, all experiments are conducted at scale, which presents a nontrivial engineering challenge for the scalability of the involved software components. We base our experiments on the network consensus from 2023-10-01 at 00:00, and simulate one hour of full network activity.
@@
Given a specific snapshot of the Tor network (consensus), the first step is to generate a trace of Tor network activity. In particular, this involves generating circuits and using them to transport data cells, simulating network communication.
Simulating the Tor network involves a trade-off between computational effort and realism. The most realistic approach uses the shadow simulator [jansen12shadow], capable of running the unmodified Tor software. However, running such simulations at scale 1:1 is infeasible. While shadow enables fine-grained simulation to investigate performance and other networking effects, we focus on the distribution of circuits and cells across the network, because these factors impact anonymity as considered by progressive pruning.
@@
We therefore adopt a more abstract simulation approach, similar to the Tor path simulator, TorPS [johnson2013usersrouted]. TorPS simulates circuit generation over time and has been used to study anonymity changes before. However, TorPS suffers from severe scalability issues. Furthermore, it does not simulate user behavior at the packet level and thus cannot be used to generate the network traces we require.
@@
We developed the Tor Flow Simulator (TorFS), an open-source simulator [^4] https://github.com/cdoepmann/torfs designed to simulate Tor user behavior at the packet level. TorFS inherits much of TorPS’s functionality. However, it is implemented in Rust with a focus on scalability and adds the missing pieces, including the modeled circuit selection rules and users’ circuit management. TorFS also supersedes TorPS’s simplistic user model, replacing it with the state-of-the-art model [DBLP:conf/ccs/JansenTH18]. That is, user behavior is now modeled as a probabilistic process (an exponential distribution) with parameters derived from real-world measurements on the Tor network. Moreover, while TorPS simulates only the generation and management of circuits, TorFS also simulates the circuits themselves to generate network traces of the individual messages (cells). For this, it again makes use of the probabilistic model from [DBLP:conf/ccs/JansenTH18]. Consequently, TorFS uses the stream model as well as the packet model from [DBLP:conf/ccs/JansenTH18], which mimic real traffic behavior as observed on the real Tor network. As a result, TorFS outputs message traces ready for analysis with ppcalc. Please note, however, that TorFS does not constitute a full-fledged network simulator like shadow or ns-3. It abstracts from networking effects in that cells entering the network are output after a constant latency at their destination.
Applying Progressive Pruning to Tor:
@@
In our evaluation, we focus on traffic that is directed from the exit relay to the client as this constitutes the majority of Tor traffic. We apply the following mapping from simulated Tor behavior to the message model from Section 3: Each cell that occurs in the simulation corresponds to a message in the model. Multiple cells that form a data stream in the simulation also constitute a single stream of messages in the sense of our model. The sender of these messages is the according simulated exit. For each stream, we define a receiver as the entity that receives all of this stream’s messages. This way, the stream-based anonymity can be evaluated independently for each stream in the network.
@@
After generating message traces of Tor, we analyze these using progressive pruning in order to calculate anonymity sets. For this task, we again use ppcalc. As laid out before, anonymity sets can be calculated in one of two possible “directions”. In our current experiment, we have a fixed (relatively small) set of exit relays that are sending data to many clients. Due to this 1:n relationship, it makes sense to calculate anonymity from the receiver perspective. That is, for each receiver (client), we calculate the set of potential senders (exits) from which their data may originate. Analyzing anonymity in the reverse direction (the set of potential receivers per sender) is not useful in this scenario due to the fact that exits are sending to many clients at the same time.
@@ 5.2 Results and Discussion
Figure 4: Anonymity sets.
Figure 5: Anonymity sets by stream length (percentiles).
[…]
@@
Our analysis is based on anonymity set sizes and their distribution across the network. Table 1 displays basic statistics about the anonymity in each simulated network. As a first observation, we can see that the anonymity set sizes are on average relatively large. In the case of the original Tor network, e.g., streams still remain with an average of more than 1,200 sender candidates (out of approximately 2,400 exits). It is noteworthy that, with regard to the simple window-based attack progressive pruning mimics, anonymity of the vast majority of streams is not trivially broken. This observation is supported by the fact that—looking again at the original Tor network—out of 26 million streams, less than 1,000 are fully de-anonymized, i.e., resulting in an anonymity set size of 1.
@@
In order to better understand how anonymity is distributed within the network, we consider CDF plots of the anonymity set sizes, displayed in Figure 5. Looking at the data from the original Tor network first, it becomes immediately apparent that anonymity is distributed very unevenly between the streams. Note that, if all streams had the same anonymity set size, the CDF plots would exhibit a single vertical line. In contrast, the simulation results show that there are many streams with very little anonymity remaining, just like there are many streams with rather strong anonymity remaining and many streams in between. While a result like this was to be expected due to the fact that Tor has no means of ensuring a certain level of anonymity, it still shows a surprisingly clear weakness of the system because users cannot reasonably expect a certain level of anonymity, as far as the stream-based anonymity degradation over time is concerned.
@@
Having seen in Section 4 that a stream’s anonymity is heavily influenced by its length, we next investigate the relationship between stream lengths and anonymity in the simulated networks. Figure 5 shows CDF plots of the anonymity sets, subdivided into quantiles based on their length (number of cells), from percentiles 0–10 (black), to 90–100 (orange). The resulting anonymity distributions clearly point out that stream length is again a decisive factor for anonymity. Long-lasting streams are massively disadvantaged and generally achieve only little anonymity. In contrast, short streams are much harder to de-anonymize and have anonymity sets that are multiple times larger. These results again support our observations from Section 4: Long streams suffer from reduced anonymity but act as cover traffic providing stronger anonymity for shorter streams.
Anonymity in Scaled Networks:
@@
We leverage our simulation and analysis workflow to evaluate not only the Tor network as is today but also variations of the network that could result from future network growth with different characteristics. Evaluating scaled Tor networks allows us to investigate the influence of network topology on anonymity. Understanding the trade-offs of different topological properties may provide important insights on how to shape and prepare future network evolution. In addition to a snapshot of today’s Tor network, we therefore carried out simulations of the Tor network being scaled by +100% vertically as well as by +100% horizontally. By “vertical” growth we mean that the bandwidth capacity of each relay is increased accordingly. Likewise, “horizontal” growth is achieved by adding new, additional relays to the existing network. The scaling methodology of torsynth is detailed in [DBLP:conf/IEEEares/DopmannT23]. For simulations, we also grow the number of active users on the network by +100% so the networks differ only in their topology, and not in the amount of available resources per user.
@@
We first consider the case of horizontal growth. Table 1 and Figure 5 show that increasing the number of relays has a strong impact on anonymity as defined by progressive pruning. In fact, with double the number of relays, the average anonymity size is also exactly doubled. This trend is intuitively reasonable because doubling the number of relays also implies doubling the number of available exits which are the potential anonymity set candidates. However, it is interesting to see that this increase of average anonymity set size appears to be exactly proportional to the number of exit relays. This is somewhat surprising because, essentially, the number of exits only directly defines the maximum anonymity set size which wouldn’t necessarily have to correspond linearly to the average anonymity set size. Even if the anonymity set sizes grow with horizontal scaling of the network, it has to be noted that the standard deviation grows massively, too. In our experiments, it also exactly doubles. We regard this as a problematic property because it means that anonymity is distributed much more unevenly among users of the network.
@@
All in all, we can conclude from the simulations of the Tor network which we carried out that there are two fundamental factors for anonymity in the network: Firstly, an increased number of relays also increases the number of (topological) candidates for the anonymity set. And secondly, the mere increase in network load also improves anonymity, albeit much less effectively.